A ring of radius 6 cm is given a charge 10 MC. how much work will be done in transporting a charge of 6 MC from its centre to a point 8 cm along its axis ?
 

Grade:12

2 Answers

Vinmuhilan
5 years ago
Workdone =6MC(V at 8cm - V at centre)
Pot. V due to a charged ring at any points 
On the axis = kq/(R^2+X^2)^0.5
Pot at centre = kq/R [ Bcoz at centre X =0]
Here X= 8cm ,q=10MC , R=6cm.
Veeresh
4 years ago
W=-q(Vfinal - V initial)
W=q(Kq/Rinitial-Kq/Rfinal)
W=140×10^-9{K(10×10^-6)/6×10^-2}-{K(10×10^-6)/8×10^-2